OpenStudy (anonymous):

FAST!!!!!!! Which of the following is equivalent to log9 36 ?

OpenStudy (lgbasallote):

if you want fast you should use a calculator...

OpenStudy (anonymous):

SORRY IDK HOW TO DO IT!

OpenStudy (kropot72):

\[\log_{9} 36=x\] \[9^{x}=36\] \[x \times \ln 9=\ln 36\] \[x=\frac{\ln 36}{\ln 9}\] Can you finish it now?